in a uniform electric field near the surface of earth, a particule having a charge of -2.0*10^-9C is acted on by a downward electric force of 3.0*10^-6N,


1. What is the magnitude and direction of the electric force exerted on a proton placed in the field?


2. What is the gravitational force on the proton?


3. What is the ratio of the electric force to the gravitational force in the case

1
Expert's answer
2021-11-22T16:15:52-0500

1) Let's first find the magnitude and direction of the electric field:


"F=qE,""E=\dfrac{F}{q}=\dfrac{3.0\times10^{-6}\ N}{-2.0\times10^{-9}\ C}=-1.5\times10^3\ \dfrac{N}{C}."

The magnitude of the electric field is "1.5\times10^3\ \dfrac{N}{C}", direction is upward.

Then, we can find the magnitude and direction of the electric force exerted on a proton placed in the field:


"F_{e,p}=q_pE=1.6\times10^{-19}\ C\times(-1.5\times10^3\ \dfrac{N}{C})=-2.4\times10^{-16}\ N."

The magnitude of the electric force exerted on a proton is "2.4\times10^{-16}\ N". The direction is upward.

2) We can find the gravitation force on the proton as follows:


"F_{g,p}=m_pg,""F_{g,p}=1.67\times10^{-27}\ kg\times9.8\ \dfrac{m}{s^2}=1.64\times10^{-26}\ N."

3) We can find the ratio of the electric force to the gravitational force as follows:


"\dfrac{F_{e,p}}{F_{g,p}}=\dfrac{2.4\times10^{-16}\ N}{1.64\times10^{-26}\ N}=1.46\times10^{10}."